OpenStudy (anonymous):

0.05x+0.25y=11 and 0.15x+0.05y=12 solve by elimination

OpenStudy (auctoratrox):

x = 70 and y = 30. multiply the entire second equation by 5 and subtract it from the first equation, then solve for x, plug in value for x into whichever equation you want to solve for y.

OpenStudy (anonymous):

first multiply the first equation by 3 so that the x term will be cancelled out then subtract the 2 equations

OpenStudy (auctoratrox):

Traile can you check your math when subtracting the first two equations?

OpenStudy (anonymous):

(0.05x+0.25y=11 )3 0.15x+0.05y=12 0.15x+0.75y=33 0.15x+0.05y=12 -0.15x-0.75y=-33 -0.7y=-21 y=30 0.15x+0.05(30)=12 x=70 ahhaahha @Auctoratrox tnx